OpenStudy (anonymous):

how do you write an equation like 13-6x+(2x-5^2)+3 in standard form???

OpenStudy (anonymous):

5^2-4x+16 i think

OpenStudy (anonymous):

5^2 =25, and 25+16 =41 so it is -4x +41 or 41-4x are good

OpenStudy (anonymous):

rewrite from largest coefiicient to least -5x^2 +2x - 6x +13+3 = -5x^2 -4x + 16